Question 9 (1 point)

Achrondroplasia is a form of dwarfism caused by a dominant allele. The homozygous dominant genotype causes death, so individuals who have the condition are all heterozygous. If a person with achondroplasia mates with a person who does not have achondroplasia, what percentage of their children would be expected to have achondroplasia

Question 9 options:

A)

0%

B)

50%

C)

75%

D)

100%

Question 10 (2 points)

This is an example of incomplete dominance.  Let H-curly hair and h-straight hair.

An individual with naturally curly hair and an individual with naturally straight hair mate. All their offspring have naturally wavy hair. If an individual with wavy hair mated with an individual with straight hair, what is the probability that their child will have curly hair?

Question 10 options:

A)

50 %

B)

25%

C)

75%

D)

0%

Question 11 (2 points)

Duchenne muscular dystropy is a sex-linked disorder (like hemophilia) characterized by a progressive loss of muscle tissue.

Neither Rudy nor Carla has Duchenne MD, but their first son does have it. If the couple has a second child, what is the probability that he or she will also have the disease? Look at all offspring possibilities, not just male or female.

Question 11 options:

A)

0 chance

B)

25%

C)

50%

D)

100%

Question 12 (1 point)

Which one of the following is an example of pleiotropy in humans?

Question 12 options:

A)

hypercholesterolemia

B)

achondroplasia

C)

Huntington's disease

D)

sickle-cell disease
Answer #1

9. Normal person = aa

Achondroplasia = Aa

Aa × aa = 1/2 Aa, 1/2 aa

Proportion of achondroplastic progenies = 1/2

Percentage of achondroplastic progenies = 1/2 × 100 = 50%

10. Straight hair = SS

Curly hair = ss

Wavy hair = Ss

Ss × SS = 1/2 SS, 1/2 Ss

Proportion of progenies with curly hair = 0

Percentage of progenies with curly hair = 0%

11. The son is suffering from DMD. It means the mother will be the carrier for this trait and the father will be normal.

Normal = XX, XdX (carrier female), XY

Affected = XdXd, XdY

XdX × XY

Proportion of affected progenies = 1/4

Percentage of affected progenies = 1/4 × 100 = 25%

Gametes X Y
Xd XdX normal XdY affected
X XX normal XY normal

12 .Hypercholesterolemia

Please rate high.